# Local Spherical Collapsing Box in Athena++: Numerical Implementation and Benchmark Tests

### v1

@@ 5 Discussion
The complete series of tests presented in §4, including the challenging diagonal wave test that induces shear flows, demonstrate the robustness and stability of our Athena++ code implementation, confirming its capability to handle complex local model simulations effectively. Athena++ has also been used for incorporating similar local models, namely the uniform expanding box for studying stellar winds (Squire et al., 2020) and the anisotropic expanding box implemented in the MHD-PIC module (Sun & Bai, 2023). These previous numerical implementations have concentrated on investigating MHD characteristics and the behavior of magnetized waves. Our study focuses on a comprehensive examination of the hydrodynamic aspects of the expanding/collapsing box. Moreover, while our code is adaptable to scenarios akin to the aforementioned expanding cases, here we focus on collapsing systems. This is motivated not only by our interest in probing protostellar collapse, but also by the fact that collapsing environments - with their intrinsic amplification of density, velocity, vorticity, and energy - are potentially more abundant in instabilities. Indeed, our observations of the growth of shear flows and wave-induced shear underline the possibility of hydrodynamic instabilities during protostellar collapse. Moreover, related previous work (Robertson & Goldreich, 2012) also presented increase of turbulent velocities (“adiabatic heating”) of contracting gas. It is thus both numerically and scientifically important to conduct a thorough investigation and validation of this local collapsing box model before delving into MHD studies.
@@
The magnetic field plays an important role in protostellar collapse, as the core must lose angular momentum during the collapse (see e.g. review of McKee & Ostriker, 2007), and magnetic braking has been considered as one of the main mechanisms to drive angular momentum transfer during core collapse and form protoplanetary disks with realistic properties (e.g. Allen et al., 2003; Price & Bate, 2007; Hennebelle & Fromang, 2008; Commerçon et al., 2011; Masson et al., 2016; Machida et al., 2016; Wurster et al., 2016; Hennebelle et al., 2020, etc.). The extension to MHD for the theoretical model of expanding/collapsing box will likely follow a similar approach to the MHD eccentric shearing box (Ogilvie & Barker, 2014). The numerical implementation, will presumably be similar to the approach described in this paper. However, magnetic fields in collapse processes may break spherical symmetry (Galli & Shu, 1993; Hennebelle, 2001; Galli et al., 2006), and thus may require further modifications of the local model. One should also be careful with the treatment of the background magnetic field gradient, in order to avoid spurious instabilities (Latter & Papaloizou, 2017).
@@
The local deviation from the spherical symmetry, such as caused by the magnetic field as discussed above, leads to the need for further generalization of this local box towards more realistic collapse profiles. A weak rotation in the global picture can be incorporated by simply using shear-periodic boundary conditions, whereas a strong rotation requires modification of the local model. Another factor to be considered is physical collapse profiles. As shown in §4.2.2, a time-dependent collapse profile can transition between different regimes, including from subsonic to supersonic flows, or from WKB to freeze-out wave regimes. It is thus necessary to study the effect of various collapse profiles, such as the typical profile of Shu (1977).
@@
Dust grains are the building blocks of planets. Recent surveys of protoplanetary disks across different stages of disk evolution (e.g. Pascucci et al., 2016; Barenfeld et al., 2016; Ansdell et al., 2018; Long et al., 2018; Cieza et al., 2019; Tychoniec et al., 2020; Tobin et al., 2020; van der Marel & Mulders, 2021) show that the available solid mass to form planets is rapidly depleted during the early stage of disk evolution, likely indicating early start of planet formation. Observational(Pagani et al., 2010; Kataoka et al., 2015; Galametz et al., 2019; Valdivia et al., 2019) and theoretical (e.g. Ormel et al., 2009; Hirashita & Yan, 2009; Guillet et al., 2020; Bate, 2022, etc.) works both suggest that dust grains are growing in collapsing protostellar cores, significantly prior to the protoplanetary disk phase. With further incorporation of dust, our local collapse simulation will be capable of resolving and investigating dust dynamics and their interactions with the gas, especially with potential instabilities and other local phenomena, which have a rich potential to concentrate dust during the collapse. Moreover, the interaction between dust and gas drag force might give rise to additional dust-related instabilities. One example, the streaming instability (Youdin & Goodman, 2005), has already been proven to be a promising avenue for planetesimal formation in protoplanetary disks, and has recently been extended to a family of instability caused by dust-gas interaction (resonance drag instability, Squire & Hopkins, 2018), which could potentially operate to concentrate dust in the context of protostellar collapse. These dust concentration mechanisms in the initial collapse phase might significantly expedite the onset of planet formation, potentially starting before the complete disk formation. This aligns with the aforementioned observational indicators of early planet formation in very young stellar systems. The implementation of dust for the current model is likely similar to the method presented in this paper. However the relative drift between the background flows of gas and dust has to be implemented in an ad-hoc way, similar to the approach of the shearing box (e.g., Bai & Stone, 2010).
@@ 6 Conclusion
We implement the local collapsing box model of LL23 into Athena++, and perform benchmark tests for the behavior of nonlinear and linear solutions in the local model. This model features a spatially uniform time-dependent Cartesian geometry, allowing (shear-)periodic boundary conditions. Key points of this study are:
  • We numerically implement the local model by modifying the Riemann solver with a rescaled anisotropic effective sound speed and an adapted energy equation, together with operator split source terms.
@@
  • Our code generally achieves convergence between first- and second-order for the diagonal wave-generated shear flow test. Further improvement in performance with adiabatic EOS is necessary for future work.
@@ Appendix A Adiabatic equation of state
The eigenvalues and eigenmatrices in the conserved variables with an adiabatic equation of state are needed to construct the fluxes of the conserved variables in the Roe’s solver. These quantities are provided hereafter
@@ Appendix C Conservation of entropy
In order to validate the conservation law of entropy in the local collapsing box, we present the results for the horizontal shear flow test runs (see §4.1.1) with adiabatic EOS, using Roe and HLLC solvers, for examining the conservation law. These runs are marked with the same labels as in Table 1.

### v2

@@ 5 Discussion
The complete series of tests presented in Sect. 4, including the challenging diagonal wave test that induces shear flows, demonstrates the robustness and stability of our Athena++ code implementation, confirming its capability to handle complex local model simulations effectively. Athena++ has also been used to incorporate similar local models; namely, the uniform expanding box for studying stellar winds (Squire et al., 2020) and the anisotropic expanding box implemented in the MHD-PIC module (Sun & Bai, 2023). These previous numerical implementations have concentrated on investigating MHD characteristics and the behavior of magnetized waves. Our study focuses on a comprehensive examination of the hydrodynamic aspects of the expanding and collapsing box. Moreover, while our code is adaptable to scenarios akin to the aforementioned expanding cases, here we focus on collapsing systems. This is motivated not only by our interest in probing protostellar collapse, but also by the fact that collapsing environments - with their intrinsic amplification of density, velocity, vorticity, and energy - are potentially more abundant in instabilities. Indeed, our observations of the growth of shear flows and wave-induced shear underline the possibility of hydrodynamic instabilities during protostellar collapse. Moreover, related previous work (Robertson & Goldreich, 2012) also presented an increase in turbulent velocities (“adiabatic heating”) of contracting gas. It is thus both numerically and scientifically important to conduct a thorough investigation and validation of this local collapsing box model before delving into MHD studies.
@@
The magnetic field plays an important role in protostellar collapse, as the core must lose angular momentum during the collapse (see e.g. review of McKee & Ostriker, 2007), and magnetic braking has been considered as one of the main mechanisms driving angular momentum transfer during core collapse and forming protoplanetary disks with realistic properties (e.g. Allen et al., 2003; Price & Bate, 2007; Hennebelle & Fromang, 2008; Commerçon et al., 2011; Masson et al., 2016; Machida et al., 2016; Wurster et al., 2016; Hennebelle et al., 2020, etc.). The extension to MHD for the theoretical model of the expanding or collapsing box will likely follow a similar approach to the MHD eccentric shearing box (Ogilvie & Barker, 2014). The numerical implementation will presumably be similar to the approach described in this paper. However, magnetic fields in collapse processes may break spherical symmetry (Galli & Shu, 1993; Hennebelle, 2001; Galli et al., 2006), and thus may require further modifications of the local model. One should also be careful with the treatment of the background magnetic field gradient, in order to avoid spurious instabilities (Latter & Papaloizou, 2017).
@@
The local deviation from the spherical symmetry, such as that caused by the magnetic field discussed above, leads to the need for further generalization of this local box toward more realistic collapse profiles. A weak rotation in the global picture can be incorporated by simply using shear-periodic boundary conditions, whereas a strong rotation requires modification of the local model. Another factor to be considered is physical collapse profiles. As is shown in Sect. 4.2.2, a time-dependent collapse profile can transition between different regimes, including from subsonic to supersonic flows, or from WKB to freeze-out wave regimes. It is thus necessary to study the effect of various collapse profiles, such as the typical profile of Shu (1977).
@@
Dust grains are the building blocks of planets. Recent surveys of protoplanetary disks across different stages of disk evolution (e.g. Pascucci et al., 2016; Barenfeld et al., 2016; Ansdell et al., 2018; Long et al., 2018; Cieza et al., 2019; Tychoniec et al., 2020; Tobin et al., 2020; van der Marel & Mulders, 2021) show that the available solid mass to form planets is rapidly depleted during the early stage of disk evolution, likely indicating an early start of planet formation. Observational(Pagani et al., 2010; Kataoka et al., 2015; Galametz et al., 2019; Valdivia et al., 2019) and theoretical (e.g. Ormel et al., 2009; Hirashita & Yan, 2009; Guillet et al., 2020; Bate, 2022, etc.) works both suggest that dust grains are growing in collapsing protostellar cores, significantly prior to the protoplanetary disk phase. With further incorporation of dust, our local collapse simulation will be capable of resolving and investigating dust dynamics and their interactions with gas, especially with potential instabilities and other local phenomena, which have a rich potential to concentrate dust during the collapse. Moreover, the interaction between dust and gas drag force might give rise to additional dust-related instabilities. One example, the streaming instability (Youdin & Goodman, 2005), has already been proven to be a promising avenue for planetesimal formation in protoplanetary disks, and has recently been extended to a family of instability caused by dust-gas interaction (resonance drag instability, Squire & Hopkins, 2018), which could potentially operate to concentrate dust in the context of protostellar collapse. These dust concentration mechanisms in the initial collapse phase might significantly expedite the onset of planet formation, potentially starting before the complete disk formation. This aligns with the aforementioned observational indicators of early planet formation in very young stellar systems. The implementation of dust for the current model is likely similar to the method presented in this paper. However, the relative drift between the background flows of gas and dust has to be implemented in an ad hoc way, similar to the approach of the shearing box (e.g., Bai10).
@@ 6 Conclusion
We implemented the local collapsing box model of LL23 into Athena++, and performed benchmark tests for the behavior of nonlinear and linear solutions in the local model. This model features a spatially uniform time-dependent Cartesian geometry, allowing (shear-)periodic boundary conditions. The key points of this study are:
  • We numerically implemented the local model by modifying the Riemann solver with a rescaled anisotropic effective sound speed and an adapted energy equation, together with operator split-source terms.
@@
  • Our code generally achieves convergence between the first and second orders for the diagonal wave-generated shear flow test. Further improvement in performance with an adiabatic EoS is necessary for future work.
@@ Appendix A Adiabatic equation of state
The eigenvalues and eigenmatrices in the conserved variables with an adiabatic EoS are needed to construct the fluxes of the conserved variables in the Roe’s solver. These quantities are provided hereafter
@@ Appendix C Conservation of entropy
In order to validate the conservation law of entropy in the local collapsing box, we present the results for the horizontal shear flow test runs (see Sect. 4.1.1) with adiabatic EoS, using Roe and HLLC solvers, for examining the conservation law. These runs are marked with the same labels as in Table 1.